An object of mass 4kg moves round a circle of radius 6m with a constant speed of 12m/s. Calculate the force towards the center.

1
Expert's answer
2022-02-28T10:41:00-0500

The centripetal force

"F=\frac{mv^2}{r}=\frac{4*12^2}{6}=96\:\rm N"